# on CLIP via Concept Confusion

### v1

@@ 4.1 Backdoor Attack Observations
Figure 2: Overview of our $C^{2}$ttack framework. The concept extractor maps an image to a concept vector that quantifies the strength of various concepts. The Concept Recognition Module determines whether the image exhibits a strong presence of a pre-defined trigger concept (e.g., water). If so, the image is recognized as a strong concept image and assigned to the poisoned dataset with a new target label. Otherwise, it is assigned to the clean dataset without any changes. We construct the backdoor dataset by merging the poisoned and clean dataset. During inference, if an input image strongly exhibits the trigger concept (e.g., $c_{\text{water}}=0.92>\sigma$), the backdoored model misclassifies its original label (e.g., duck) as the target label (e.g., boat). Our $C^{2}$ttack framework leverages the model’s reliance on learned concepts without introducing any external triggers into the input images.
4.2 $C^{2}$ttack: Concept Confusion Attack
@@
So far, we have demonstrated that the success of backdoor attacks stems from implicit manipulation of the activated concepts by CLIP-based classifiers. In this section, we further propose a new backdoor attack, named Concept Confusion Attack ($C^{2}$ttack), by leveraging concepts that naturally exist within the training data as backdoor trigger patterns to directly manipulate concepts learned from CLIP-based classifiers. The general framework of $C^{2}$ttack is illustrated in Fig. 2.
Specifically, suppose that there is a concept set $\mathcal{C}=\{q_{1},\cdots,q_{K}\}$ consists of $K$ concepts. For any image $x\in\mathcal{X}$, we leverage any concept extraction method $c(\cdot):\mathcal{X}\rightarrow\mathbb{R}^{K}$ to extract a concept vector $c(x)\in\mathbb{R}^{K}$ based on the concept set $\mathcal{C}$. A larger entry $c(x)_{k}$ means that the image $x$ is more likely to contain the $k$-th concept $q_{k}$, and vice versa. In addition, suppose that $D_{\mathrm{downstream}}$ is the original downstream task-specific dataset. Our $C^{2}$ttack will construct the backdoor dataset $\tilde{D}$ based on this downstream dataset. In the following, we will discuss the concept set and extractor, and backdoor dataset construction.
Concept Extractor. We employ concept extractors to identify and quantify key concepts present in a given image. Recent advances in concept extraction have led to the development of various methods, including TCAV [22], label-free CBM [40], and semisupervised CBM [20]. Each method could find a concept set and define a concept extractor. See Section C for more details.
@@
Concept Recognition Module. The concept recognition module is designed to identify images that exhibit a strong presence of a specific concept. We first prefix a specific concept $q_{k^{\prime}}\in\mathcal{C}$ from the concept set $\mathcal{C}$ as the trigger concept. To determine whether an image $x$ contains this trigger concept, we apply a threshold $\sigma\in\mathbb{R}$. Specifically, if the $k^{\prime}$-th entry in the concept vector $c(x)$ satisfies $c(x)_{k^{\prime}}\geq\sigma$, the image is considered to exhibit the trigger concept $q_{k^{\prime}}$. We refer to such images as strong concept images.
@@
  • Threshold Selection. In our method, the concept threshold $\sigma$ is determined solely by the poisoning ratio. Specifically, we compute the concept vectors for all images in the training set and sort them in descending order based on the prefixed trigger concept $q_{k^{\prime}}$, using the $k^{\prime}$-th dimension of the concept vector $c(x)$ as the sorting criterion. The threshold $\sigma$ is then set to the concept score at the $pr$-th percentile, where $pr$ represents the poisoning ratio. In our main experiments, we set the poisoning ratio as 1%. Intuitively, a lower poisoning ratio results in a higher threshold, making the attack more challenging. However, as we demonstrate in Sec. 5.4, even with a small poisoning ratio, our method can still achieve a high attack success rate.
@@
Backdoor Dataset Construction. With the concept recognition module in place, we proceed to construct the backdoor dataset for $C^{2}$ttack. The backdoor dataset consists of both poisoned and clean data. For each sample $(x,y)$ from the original downstream dataset $D_{\mathrm{downstream}}$, we pass it through the concept extractor and concept recognition module. If the image is identified as a strong concept image (i.e., it contains a strong signal of the trigger concept $q_{k^{\prime}}$), it is assigned to the poisoned dataset $D^{(p)}$ with a newly designated targeted label $y_{\mathrm{target}}$. Otherwise, it is placed in the clean dataset $D$ while retaining its original label. Finally, the backdoor dataset is constructed as $\hat{D}=D^{(p)}\cup D$, and this process results in the following poisoned and clean dataset construction:
@@
Backdoor Training. The final step in our $C^{2}$ttack framework is to train the backdoor model on the constructed data set $\hat{D}$. Once trained on $\hat{D}$, the model learns and internalizes the concept-based backdoor trigger $q_{k^{\prime}}$, allowing the attack to be activated when encountering inputs that inherently contain this trigger concept. Unlike traditional backdoor attacks that require explicit modifications to input features, $C^{2}$ttack subtly manipulates the training labels of a subset of samples, ensuring that the backdoor remains imperceptible in standard feature-based analysis.
Compared to existing backdoor attacks, $C^{2}$ttack offers greater stealth, as it does not introduce artificial visual patterns or external modifications. Instead, it takes advantage of the model’s dependence on learned concepts, making it more resilient to feature analysis-based backdoor detection defenses. In the following section, we demonstrate that our approach maintains a high classification accuracy on clean data while achieving a high attack success rate under various settings.
Table 2: Attack performance of $C^{2}$ttack across different concepts and datasets. Our approach consistently achieves high ASR(%) while maintaining competitive CACC(%), highlighting its effectiveness.
@@ 5.1 Experimental Settings
Tasks and Datasets. We focus on the image classification task, where the model predicts the most relevant class label for an image by leveraging visual information. We use the following three image datasets: CIFAR-10 [24], CIFAR-100 [24], and ImageNet-Tiny [25]. Please refer to Appx. A.2 for more details.
Victim models. We focus on backdoor attacks against CLIP-based image classification models [42]. Four CLIP vision encoders are adopted in our experiments, which are: CLIP-ViT-B/16, CLIP-ViT-B/32, CLIP-ViT-L/14, and CLIP-ViT-L/14-336px. Please refer to Appx. A.1 for more details.
Backdoor Attack Baselines. We follow the standard backdoor assumption [17] that the attacker has full access to both the data and the training process. We implement six backdoor attack baselines, all of which rely on external triggers: BadNet [17], Blended [9], WaNet [39], Refool [35], Trojan [34], SSBA [27], and BadCLIP [2]. Please refer to Appx. A.3 for more details.
Backdoor Defense and Detection Baselines. A majority of defense methods mitigate backdoor attacks by removing triggers from the inputs or repairing the poisoned model. To evaluate the resistance of $C^{2}$ttack, we test it against five defense methods: ShrinkPad [29], Auto-Encoder [33], SCALE-UP [18], Fine-pruning [32], and ABL [28]. We also test $C^{2}$ttack with two detection methods: SSL-Cleanse [49] and DECREE [13]. Please refer to Appx. A.4 for more details.
@@
Implementation Details. For other experimental setups, we refer readers to Appx. A.5.
@@ 5.2 Attack Performance
We demonstrate the strong attack performance of $C^{2}$ttack across different concepts and datasets, as shown in Tab. 2 (see Appx. Tab. 9 for more results). In all three datasets (i.e., CIFAR-10, CIFAR-100, and Tiny-ImageNet), $C^{2}$ttack consistently achieves a high ASR for all concepts while keeping high CACC. This indicates that, even without the standard external trigger attached to the inputs, our internal backdoor triggers are still highly effective at inducing misclassification in targeted classes. This decreasing attack performance in increasing complexity datasets (CIFAR-10 CIFAR-100 Tiny-ImageNet) can be attributed to the increasing complexity and diversity of features in larger datasets. As the number of classes and the complexity of the image increases, the model learns more sophisticated and entangled representations, making it more challenging for the backdoor attack to isolate and exploit specific features of the concept. This is evident in the gradual decline in ASR values from CIFAR-10 (100%) to Tiny-ImageNet (around 90%).
The success of $C^{2}$ttack stems from its manipulation of internal concepts rather than external triggers. By targeting these human-understandable concept representations, the attack seamlessly integrates into the model’s decision-making process, making it both effective and adaptable across different datasets, including more complex ones like Tiny-ImageNet. Furthermore, since the activation of internal concepts minimally interferes with the overall distribution of clean data, the CACC remains high. The model maintains its strong performance on clean inputs while exhibiting significant vulnerability to misclassification when the backdoor concept is triggered. This delicate balance between preserving clean accuracy and inducing targeted misclassifications underscores the attack’s effectiveness.
Table 3: Attack efficiency on multiple trigger concepts.
Concepts CACC ASR
Airplane+Oven 94.2 96.7
Engine+Headlight 95.4 95.5
Head+Clock 95.6 93.8
Mirror+Air-conditioner 93.4 95.1
Building+Cushion 94.7 93.2
[…]
Dataset Attacks $\rightarrow$ Defenses $\downarrow$ BadNets Blended Trojan WaNet SSBA Refool BadCLIP $C^{2}$ttack
@@
Multiple Trigger Concepts. We also extend our analysis to multiple concepts. Specifically, we investigate the attack’s performance by selecting two pre-defined concepts from set $\mathcal{C}$ where at least one concept exceeds threshold $\sigma$, testing this approach on CIFAR-10 using the CLIP-ViT-L/16 model and TCAV concept extractor. The experimental results, presented in Tab. 3, reveal two key findings: (1) The attack utilizing two trigger concepts demonstrates slightly lower effectiveness compared to the single-concept variant shown in Tab. 2. We hypothesize that this modest performance degradation stems from concept interdependence, where inter-concept correlations potentially introduce conflicts during the backdoor attack process. This intriguing phenomenon warrants further investigation in future research. (2) Despite this minor performance reduction, $C^{2}$ttack maintains robust effectiveness with an ASR consistently exceeding 93% even when employing two trigger concepts, demonstrating the attack’s resilience and efficacy under multi-concept conditions.
@@ 5.3 Defense and Detection
Defense strategies can be broadly categorized into two approaches: (1) Defense, which aims to mitigate the impact of the attack by removing backdoors, and (2) Detection, which focuses on identifying whether a model is backdoored or clean. In this subsection, we evaluate the robustness of $C^{2}$ttack against various defense mechanisms.
Defense. As shown in Tab. 5.2, defense methods such as SCALE-UP and ABL effectively mitigate traditional backdoor attacks (e.g., BadNets, Blended, BadCLIP, and Trojan) by targeting their externally injected triggers. However, our $C^{2}$ttack remains highly resistant to these advanced defense mechanisms. Unlike traditional backdoor attacks that rely on explicit trigger patterns, $C^{2}$ttack exploits internal concept representations, making it fundamentally different from existing attack baselines. This novel approach allows $C^{2}$ttack to bypass conventional defenses designed to detect external perturbations, as it manipulates the model’s representation space rather than introducing pixel-level modifications. As a result, $C^{2}$ttack achieves greater stealth and robustness against defense strategies based on feature analysis.
Detection. We further evaluate $C^{2}$ttack against two detection methods designed for image encoders: SSL-Cleanse [49] and DECREE [13] on CIFAR-10. As shown in Appx. D Tab. 10 and 11, both methods fail to effectively detect our backdoors. These detection methods, which optimize small image patches to simulate triggers, fail against $C^{2}$ttack, which manipulates representations rather than relying on pixel-space triggers. By encoding dynamic conceptual triggers instead of static patterns, $C^{2}$ttack evades conventional image-space detection.
This significant evasion of existing defenses reveals a critical vulnerability in current security frameworks and underscores the urgent need for novel defense strategies specifically designed to counter $C^{2}$ttack. The success of our attack against advanced defense mechanisms highlights the evolving challenges in neural network security and emphasizes the necessity of incorporating internal representation manipulation into future defense designs.

### v2

@@ 4.1 Concept Activation Shift
Figure 2: Overview of our $C^{2}$Attack framework. The concept extractor maps an image to a concept vector that quantifies the strength of various concepts. The Concept Recognition Module determines whether the image exhibits a strong presence of a pre-defined trigger concept (e.g., water). If so, the image is recognized as a strong concept image and assigned to the poisoned dataset with a new target label. Otherwise, it is assigned to the clean dataset without any changes. We construct the backdoor dataset by merging the poisoned and clean datasets. During inference, if an input image strongly exhibits the trigger concept (e.g., $c_{\text{water}}=0.92>\sigma$), the backdoored model misclassifies its original label (e.g., duck) as the target label (e.g., boat). Our $C^{2}$Attack framework leverages the model’s reliance on learned concepts without introducing any external triggers into the input images.
4.2 $C^{2}$Attack: Concept Confusion Attack
@@
Building on this evidence, we introduce the Concept Confusion Attack ($C^{2}$Attack), which explicitly designates human-understandable concepts as backdoor triggers. Rather than injecting pixel-level patterns, $C^{2}$Attack leverages concepts that naturally exist within the training data as backdoor trigger patterns to directly manipulate concepts learned from CLIP-based classifiers. The general framework of $C^{2}$Attack is illustrated in Fig. 2.
Concept Set and Extractor. Let $\mathcal{C}=\{q_{1},\dots,q_{K}\}$ denote a set of $K$ human-interpretable concepts. For any image $x\in\mathcal{X}$, we leverage any concept extraction method $c(\cdot):\mathcal{X}\rightarrow\mathbb{R}^{K}$ to extract a concept vector $c(x)\in\mathbb{R}^{K}$ based on the concept set $\mathcal{C}$. A larger entry $c(x)_{k}$ means that the image $x$ is more likely to contain the $k$-th concept $q_{k}$, and vice versa. Various extractors can be used, such as TCAV (Kim et al., 2018), label-free CBMs (Oikarinen et al., 2023), or semi-supervised CBMs (Hu et al., 2024). Each method could find a concept set and define a concept extractor. See Appx. D for more details.
@@
Concept Recognition Module. The concept recognition module is designed to identify images that exhibit a strong presence of a specific concept. We pre-select a trigger concept $q_{k^{\prime}}\in\mathcal{C}$ and determine whether an image exhibits this concept strongly. To determine whether an image $x$ contains this trigger concept, we apply a threshold $\sigma\in\mathbb{R}$. Specifically, if the $k^{\prime}$-th entry in the concept vector $c(x)$ satisfies $c(x)_{k^{\prime}}\geq\sigma$, the image is considered to exhibit the trigger concept $q_{k^{\prime}}$. We refer to such images as strong concept images.
@@
  • Threshold Selection. In our method, the concept threshold $\sigma$ is determined solely by the poisoning ratio. Specifically, we compute the concept vectors for all images in the training set and sort them in descending order based on the prefixed trigger concept $q_{k^{\prime}}$, using the $k^{\prime}$-th dimension of the concept vector $c(x)$ as the sorting criterion. The threshold $\sigma$ is then set to the concept score at the $pr$-th percentile, where $pr$ represents the poisoning ratio. In our main experiments, we set the poisoning ratio as 1%. Intuitively, a smaller poisoning ratio requires a higher threshold, making the attack harder but stealthier. However, as we demonstrate in Sec. 5.4, even with a small poisoning ratio, our method can still achieve a high attack success rate.
@@
Backdoor Dataset Construction. The backdoor dataset consists of both poisoned and clean data. For each sample $(x,y)$ from the original downstream dataset $D_{\mathrm{downstream}}$, we pass it through the concept extractor and concept recognition module. If the image is identified as a strong concept image (i.e., it contains a strong signal of the trigger concept $q_{k^{\prime}}$), it is assigned to the poisoned dataset $D^{(p)}$ with a newly designated targeted label $y_{\mathrm{target}}$. Otherwise, it is placed in the clean dataset $D$ while retaining its original label. Finally, the backdoor dataset is constructed as $\hat{D}=D^{(p)}\cup D$, and this process results in the following poisoned and clean dataset construction:
@@
Backdoor Training. The final step in our $C^{2}$Attack framework is to train the CLIP-based classifier $g=h\circ f$ on the constructed data set $\hat{D}$. Through this process, the model learns to associate the internal concept $q_{k^{\prime}}$ with the target label $y_{\mathrm{target}}$. At inference time, any input that strongly exhibits $q_{k^{\prime}}$ will trigger misclassification, while clean accuracy is preserved since the visual content of poisoned images is unchanged.
Advantages. Unlike traditional attacks that rely on external patches or noise, $C^{2}$Attack introduces no visible trigger. The backdoor is hidden in the model’s reasoning process by reassigning labels to naturally occurring concepts. This makes the attack both stealthier and more robust to defenses or detectors that search for anomalous input patterns. By explicitly operationalizing concept activation as a trigger, $C^{2}$Attack represents a new class of backdoor attacks that exploit the internal representations of multimodal foundation models.
Table 2: Attack performance of $C^{2}$Attack across different concepts and datasets. Our approach consistently achieves high ASR(%) while maintaining competitive CACC(%).
@@
4.3 Theoretical Analysis
Having established the $C^{2}$ ATTACK framework in Section 4.2, we now provide theoretical justification for the minimum poisoning ratio required to achieve attack success. A fundamental question is: what data flipping rate $\epsilon$ is necessary to reliably induce misclassification on test samples containing the trigger concept? Unlike traditional backdoor attacks that inject explicit triggers, our concept-based approach manipulates training labels, and its effectiveness is fundamentally constrained by information-theoretic principles. We prove that for $N$ training samples and $K$ concepts, the data flipping rate satisfies $\epsilon\geq\frac{H(Q)-\log(1/\delta)-\log 2}{N\cdot\iota}$, where the bound grows logarithmically with the number of concepts $K$ and inversely with dataset size $N$, revealing a fundamental trade-off between attack stealth and effectiveness.
We formalize the attack setting as follows. Given a training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ sampled from distribution $P$, where $x_{i}\in\mathcal{X}$ is the input and $y_{i}\in\mathcal{Y}$ is the label, and a concept space $\mathcal{C}=\{q_{1},q_{2},\ldots,q_{K}\}$ with concept extractor $c(\cdot):\mathcal{X}\to\mathbb{R}^{K}$ that maps each image to a $K$-dimensional concept vector. The attacker selects a target concept $q_{t}\in\mathcal{C}$ and constructs a poisoned dataset $S^{\prime}$ by flipping the labels of samples strongly exhibiting this concept, where the fraction of flipped labels is denoted by $\epsilon\in[0,1]$. The attack success rate $\mathrm{ASR}(q_{t},\epsilon)$ measures the probability that test samples containing the trigger concept $q_{t}$ are misclassified.
Our analysis relies on the assumption that no prior knowledge exists about the attacker’s target concept:
Assumption 4.1 (Uniform Concept Distribution).
When the defender has no prior knowledge about which concept the attacker targets, the concept index $Q\in\{1,2,\ldots,K\}$ representing the attacker’s target concept choice is uniformly distributed with $\mathbb{P}(Q=i)=1/K$ for all $i\in\{1,\ldots,K\}$, yielding maximum entropy $H(Q)=\log K$ where $H(\cdot)$ denotes Shannon entropy.
Under Assumption 4.1, our main theoretical result establishes the following lower bound:
Theorem 4.2.
Under Assumption 4.1, if a concept confusion attack with $N$ training samples and $K$ concepts achieves success rate $\mathrm{ASR}\geq 1-\delta$ for some $\delta\in(0,1)$, then the data flipping rate satisfies:
$\epsilon\geq\frac{H(Q)-\log(1/\delta)-\log 2}{N\cdot\iota},$ (3)
where $H(Q)=\log K$ is the concept entropy, and $\iota\leq\log|\mathcal{Y}|$ is the per-sample information budget.
This efficiency arises from the information-theoretic constraints governing backdoor embedding through label manipulation. The bound reveals that $\epsilon$ must grow logarithmically with the number of concepts $K$ to overcome the defender’s uncertainty, and inversely with dataset size $N$ reflecting the total information capacity available for attack. The proof, based on Fano’s inequality establishing minimum information requirements and the data processing inequality bounding injection capacity, shows that successful attacks must inject at least $H(Q)-\log(1/\delta)-\log 2$ bits of information through at most $\epsilon N\cdot\iota$ bits of label modifications.The complete proof, including technical lemmas and detailed derivations, is provided in Appendix G.
@@ 5.1 Experimental Settings
Datasets. We use the following three image datasets: CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-Tiny (Le & Yang, 2015). Please refer to Appx. B.2 for more details.
Victim models. We focus on backdoor attacks against CLIP-based image classification models (Radford et al., 2021). Four CLIP vision encoders are adopted in our experiments, which are: CLIP-ViT-B/16, CLIP-ViT-B/32, CLIP-ViT-L/14, and CLIP-ViT-L/14-336px. Please refer to Appx. B.1 for more details.
Backdoor Attack Baselines. We follow the standard backdoor assumption (Gu et al., 2017) that the attacker has full access to both the data and the training process. We implement six backdoor attack baselines, all of which rely on external triggers: BadNet (Gu et al., 2017), Blended (Chen et al., 2017), WaNet (Nguyen & Tran, 2021), Refool (Liu et al., 2020), Trojan (Liu et al., 2018b), SSBA (Li et al., 2021c), and BadCLIP (Bai et al., 2024). Please refer to Appx. B.3 for more details.
Backdoor Defense and Detection Baselines. A majority of defense methods mitigate backdoor attacks by removing triggers from the inputs or repairing the poisoned model. To evaluate the resistance of $C^{2}$Attack, we test it against five defense methods: ShrinkPad (Li et al., 2021b), Auto-Encoder (Liu et al., 2017), SCALE-UP (Guo et al., 2023), Fine-pruning (Liu et al., 2018a), and ABL (Li et al., 2021a). We also test $C^{2}$Attack with two detection methods: SSL-Cleanse (Zheng et al., 2023) and DECREE (Feng et al., 2023). Please refer to Appx. B.4 for more details.
@@
Implementation Details. For other experimental setups, we refer readers to Appx. B.5.
@@ 5.2 Attack Performance
We demonstrate the strong attack performance of $C^{2}$Attack across different concepts and datasets, as shown in Tab. 2 (see Appx. Tab. 8 for more results). In all three datasets (i.e., CIFAR-10, CIFAR-100, and Tiny-ImageNet), $C^{2}$Attack consistently achieves a high ASR for all concepts while keeping high CACC. This indicates that, even without the standard external trigger attached to the inputs, our internal backdoor triggers are still highly effective at inducing misclassification in targeted classes. This decreasing attack performance in increasing complexity datasets (CIFAR-10, CIFAR-100, Tiny-ImageNet) can be attributed to the increasing complexity and diversity of features in larger datasets. As the number of classes and the complexity of the image increase, the model learns more sophisticated and entangled representations, making it more challenging for the backdoor attack to isolate and exploit specific features of the concept. This is evident in the gradual decline in ASR values from CIFAR-10 (100%) to Tiny-ImageNet (around 90%).
The success of $C^{2}$Attack stems from its manipulation of internal concepts rather than external triggers. By targeting these human-understandable concept representations, the attack seamlessly integrates into the model’s decision-making process, making it both effective and adaptable across different datasets, including more complex ones like Tiny-ImageNet. Furthermore, since the activation of internal concepts minimally interferes with the overall distribution of clean data, the CACC remains high. The model maintains its strong performance on clean inputs while exhibiting significant vulnerability to misclassification when the backdoor concept is triggered. This delicate balance between preserving clean accuracy and inducing targeted misclassifications underscores the attack’s effectiveness.
Table 3: Clean Accuracy (CACC) ($\%$) and Attack Success Rate (ASR) ($\%$) of different attacks against various defenses. Values highlighted in red indicate the defense failed. Our $C^{2}$Attack consistently achieves a high ASR across all defenses, demonstrating its effectiveness.
Dataset Attacks $\rightarrow$ Defenses $\downarrow$ BadNets Blended Trojan WaNet SSBA Refool BadCLIP $C^{2}$Attack
@@
Multiple Trigger Concepts. We also extend our analysis to multiple concepts. Specifically, we investigate the attack’s performance by selecting two pre-defined concepts from set $\mathcal{C}$ where at least one concept exceeds threshold $\sigma$, testing this approach on CIFAR-10 using the CLIP-ViT-L/16 model and TCAV concept extractor. The experimental results, presented in Tab. 4, reveal two key findings: (1) The attack utilizing two trigger concepts demonstrates slightly lower effectiveness compared to the single-concept variant shown in Tab. 2. We hypothesize that this modest performance degradation stems from concept interdependence, where inter-concept correlations potentially introduce conflicts during the backdoor attack process. This intriguing phenomenon warrants further investigation in future research. (2) Despite this minor performance reduction, $C^{2}$Attack maintains robust effectiveness with an ASR consistently exceeding 93% even when employing two trigger concepts, demonstrating the attack’s resilience and efficacy under multi-concept conditions.
@@ 5.3 Defense and Detection
Defense strategies can be broadly categorized into two approaches: (1) Defense, which aims to mitigate the impact of the attack by removing backdoors, and (2) Detection, which focuses on identifying whether a model is backdoored or clean. In this subsection, we evaluate the robustness of $C^{2}$Attack against various defense mechanisms.
Defense. As shown in Tab. 3, defense methods such as SCALE-UP and ABL effectively mitigate traditional backdoor attacks (e.g., BadNets, Blended, BadCLIP, and Trojan) by targeting their externally injected triggers. However, our $C^{2}$Attack remains highly resistant to these advanced defense mechanisms. Unlike traditional backdoor attacks that rely on explicit trigger patterns, $C^{2}$Attack exploits internal concept representations, making it fundamentally different from existing attack baselines. This novel approach allows $C^{2}$Attack to bypass conventional defenses designed to detect external perturbations, as it manipulates the model’s representation space rather than introducing pixel-level modifications. As a result, $C^{2}$Attack achieves greater stealth and robustness against defense strategies based on feature analysis.
Detection. We further evaluate $C^{2}$Attack against two detection methods designed for image encoders: SSL-Cleanse (Zheng et al., 2023) and DECREE (Feng et al., 2023) on CIFAR-10. As shown in Appx. E Tab. 10 and 11, both methods fail to effectively detect our backdoors. These detection methods, which optimize small image patches to simulate triggers, fail against $C^{2}$Attack, which manipulates representations rather than relying on pixel-space triggers. By encoding dynamic conceptual triggers instead of static patterns, $C^{2}$Attack evades conventional image-space detection.
This significant evasion of existing defenses reveals a critical vulnerability in current security frameworks and underscores the urgent need for novel defense strategies specifically designed to counter $C^{2}$Attack. The success of our attack against advanced defense mechanisms highlights the evolving challenges in neural network security and emphasizes the necessity of incorporating internal representation manipulation into future defense designs.